OpenStudy (anonymous):

Please help will medal and fan! The function f(x) = x^3 + 4 is transformed to give a new function, g(x). The graph of g(x) [red] shows a horizontal shrink compared to the graph of f(x) [blue]. What is the function g(x)?

OpenStudy (anonymous):

g(x) = 2(x^3 + 4) g(x)=½ (x^3+4) g(x) = 2x^3 + 4 g(x)=½ x^3+4

OpenStudy (anonymous):

Please help nobody can figure out the correct answer

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

It's either c or d ibthink